OpenStudy (anonymous):

Two executives fly towards each other for a meeting. The first jet goes 320 mph, while the second goes 375 mph. If they started from opposite coasts 3100 miles apart, how long will it take them to meet? Round to the nearest tenth if necessary

OpenStudy (anonymous):

someone please help. D:

OpenStudy (ujjwal):

let they meet at a distance x from first jet. so x=320* t 3100-x =37*t where t is time taken for them to meet from above two relations find x that is your answer.

OpenStudy (ujjwal):

*375

OpenStudy (ujjwal):

I meant 3100-x = 375*t

OpenStudy (anonymous):

it's a decimal answer :o idk.

OpenStudy (anonymous):

9.7 hours 8.3 hours 2.2 hours 4.5 hours

OpenStudy (anonymous):

one of those ^ ^ ^

OpenStudy (anonymous):

:o

OpenStudy (ujjwal):

sorry, you don't need to find x you need to find t and it comes out to be4.5

OpenStudy (anonymous):

Thank's.:D

OpenStudy (anonymous):

Let the first jet cover x miles in t hours Then in same time the second jet will have covered 3100-x miles in same time t hours Now time = distance/speed time for first jet t = x/320 ---------------------- (I) time for second jet t = (3100-x)/375 -------------- (II) since t is same for both, we can equate the two x 3100-x ----- = --------- 320 375 Cross-multiplying 375x = 320(3100-x) 375x = 992000 - 320x 375x + 320x = 992000 695x = 992000 x= 992000/695 x = 1427.338 substituting in (I) t = 1427.338/320 t= 4.46 so ans is 4.5 hours